OpenStudy (anonymous):

This relation is already written in ordered pairs. The only thing you have to do is place the ordered pairs inside braces { } and name the relation! R = {(2, 3), (1, 2), (4, 2), (2, −1), (−2, 1)} EXPLAIN" dont give me the answer EXPLAIN. please help

OpenStudy (anonymous):

@ganeshie8 @jim_thompson5910

OpenStudy (anonymous):

@chrisdbest @mathslover

OpenStudy (anonymous):

i going to take a exam AND i need to understand

jimthompson5910 (jim_thompson5910):

I don't understand the question fully. What do they want exactly? Is that all it says?

OpenStudy (anonymous):

look

OpenStudy (anonymous):

A relation simply describes a relationship and pairs input values with output values. You can represent a relation in several ways. The snacks were represented as ordered pairs, like (3, pretzels). Let's look at a relation with pairs of numbers, instead of a number and a snack. Select each tab to see four ways this relation can be written. (2, 3), (1, 2), (4, 2), (2, −1), (−2, 1)

jimthompson5910 (jim_thompson5910):

so it looks like you already wrote the answer

OpenStudy (anonymous):

its just practice, but i need to understand the problem, because i going to take the exam

jimthompson5910 (jim_thompson5910):

Is this practice applet providing the steps and answer?

OpenStudy (anonymous):

never mind, explain me this

jimthompson5910 (jim_thompson5910):

can you post a full screenshot of the entire problem?

OpenStudy (anonymous):

Find the domain and range of the relation shown here in graph form.

OpenStudy (anonymous):

can you explain me this?

jimthompson5910 (jim_thompson5910):

domain = set of x values range = set of y values

jimthompson5910 (jim_thompson5910):

so you just list the x coordinates of each point to get the domain

OpenStudy (anonymous):

(3,4) (5, 8)?

jimthompson5910 (jim_thompson5910):

what are the x coordinates

jimthompson5910 (jim_thompson5910):

hint: each point is of the form (x,y)

OpenStudy (anonymous):

{−3, −1, 1, 5}?

jimthompson5910 (jim_thompson5910):

very good, that is your domain

jimthompson5910 (jim_thompson5910):

the range is similar, just with the y values

OpenStudy (anonymous):

thank you, can you explain me with more questions? plz

jimthompson5910 (jim_thompson5910):

sure I'll help with one more

OpenStudy (anonymous):

Identify the slope and y-intercept of the equation. y − 4 = −3x.

jimthompson5910 (jim_thompson5910):

add 4 to both sides to get y = ???

OpenStudy (anonymous):

y − 4 = − 3x +4 +4 y = − 3x + 4

jimthompson5910 (jim_thompson5910):

very nice

OpenStudy (anonymous):

im right?

jimthompson5910 (jim_thompson5910):

compare y = -3x + 4 to y = mx+b you'll find that m = -3 and b = 4 m is the slope b is the y-intercept

jimthompson5910 (jim_thompson5910):

since m = -3, the slope is -3 since b = 4, the y-intercept is 4. This is the point (0,4)

OpenStudy (anonymous):

yea, thank you for explain me

jimthompson5910 (jim_thompson5910):

you're welcome